OpenStudy (anonymous):

0.9(2x+8)=20-(x+5) ?

OpenStudy (anonymous):

1. distribute 2.simplify

OpenStudy (anonymous):

0.9(2x)+0.9(8)= 20-x-5

OpenStudy (anonymous):

1.8x+7.2=15-x

OpenStudy (anonymous):

add (1)x to both sides and subtract 7.2 from both sides

OpenStudy (anonymous):

2.8x=7.8

OpenStudy (anonymous):

tyy

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

um...no that's not right..- - sorry

OpenStudy (anonymous):

39/14

OpenStudy (anonymous):

that's the answer